OpenStudy (anonymous):

Rational Expression: 3/4d - 1/6 = 4/8d + 1/2

OpenStudy (anonymous):

Is this written as: \[\frac{ 3 }{ 4d } - \frac{ 1 }{ 6 } = \frac{ 4 }{ 8d } + \frac{ 1 }{ 2 }\] ?

OpenStudy (anonymous):

You can start by creating a common denominator for both sides

jhonyy9 (jhonyy9):

i think this wan being 3 1 4 1 3 1 --d - -- = --d + -- add to both sides - --d - --- 4 6 8 2 4 2 3 3 1 1 4 3 1 1 --d - --d - -- - -- = --d - --d - -- + -- 4 4 2 6 8 4 2 2 1 1 1 3 - -- - -- = --d - --d 2 6 2 4 (-3-1) (2d-3d) ------ = ------- 6 4 -4 -d --- = --- 6 4 -4*4=-d*6 -16=-6d 8=3d d= 8/3